# Antifragility of RIS-assisted Communication Systems under Jamming Attacks

### v1

@@ Abstract
Antifragility of communication systems is defined as measure of benefits gained from the adverse events and variability of its environment. In this paper, we introduce the notion of antifragility in Reconfigurable Intelligent Surface (RIS) assisted communication systems affected by a jamming attack. We analyzed the antifragility of the two hop systems, where the wireless path contains source node, RIS, destination node, and a eavesdropping/jamming node. We propose and analyze the antifragility performance for several jamming models, such as Digital Radio Frequency Memory (DRFM) and phase and amplitude shifting. Our paper shows that antifragility throughput can indeed be achieved under certain power thresholds and for various jamming models. In particular, high jamming power combined with low baseline data rates yields an antifragile gain factor of approximately five times. The results confirm that reconfigurable intelligent surfaces, when coupled with an antifragile design philosophy, can convert hostile interference from a liability into a throughput gain.
@@ I Introduction
Conventional resilient communication seeks to prevent or mitigate the effects of disruptions through hostile conditions, or attacks. Whereas resilience aims at preserving or restoring the existing functionality, a new paradigm, referred to as antifragility, goes a step further by leveraging uncertainty and adverse conditions for performance gains. According to its original principles first presented in [1], antifragile systems thrive under randomness, by improving their operational metrics when faced with challenges that would simply degrade or disrupt other systems. The first work on application of antifragility in communication systems by [2] showed that network throughput can greatly improve when subjected to adversarial jamming attacks. Despite the growing interest in designing and operating future mobile networks reliably, antifragility remains a relatively under-explored concept in communications and networking.
We are interested in the specific research question which is whether one new technology carries to potential for antifragile performance gains, i.e., Reconfigurable Intelligent Surfaces (RIS). Aside from being specific, our research question is also relevant, since RIS, by employing a two-dimensional array of reflecting elements, where the parameter, such as amplitude and phase, can dynamically be tuned, is known to effectively mitigate interference and extend coverage [3]. It is this capability especially that carries potential to achieving antifragile gains under jamming attacks. Analyzing RIS-assisted systems under the jamming scenarios is therefore fundamental to assessing their adaptability and to developing strategies that leverage deliberate interference for performance gains.
In this paper, we study analytically whether RIS-assisted communication systems can achieve antifragile performance gains under jamming attack. We focus on a two-hop communication system consisting of source node, RIS, and destination node, with an eavesdropping or jamming node acting as an adversarial entity. We subject this system to various jamming models, including Digital Radio Frequency Memory (DRFM) and phase and amplitude shifting. Our analysis reveals that by leveraging the reflections of the RIS, communication systems subjected to high jamming power can, in fact, experience improved throughput under certain conditions, demonstrating a remarkably antifragile behavior. Specifically, when the system baseline data rate is low and adversarial power is high, we observe a notable antifragile gain factor of up to 5×5$\times$5 ×. Although the gain factor is smaller when baseline data rates are larger, it still represents a notable improvement under strong jamming. By addressing a range of jamming techniques and analyzing the role of RIS in achieving antifragility, this work provides a foundation for developing more generally adaptive, self-improving wireless networks.
The rest of the paper is organized as follows. Section II presents the related work. Section III provides the system model and Section IV the antifragility scheme design. Section V discusses numerical results. Section VI concludes the paper.
@@ II Related Work
Paper [4] presented a method to carefully design frequency-shift keying (FSK) waveforms in order to exploit reactive jammers, effectively forcing an attacker, i.e., the jammer node, to act as an unintentional relay and thereby enhance the data rates of the legitimate users. Building on this fundamental idea, the early work in [2] studied the antifragile wireless communications showing the antifragile throughput gain under reactive jamming scenarios. More recently, [5], basing its antifragility scheme on paper [4] focused on reducing the outage probability in multi-relay cognitive networks, however without consideration of the jamming model type, signal separation and spoofing signal creation.
Our study is the first to explore antifragility achieved in wireless communications with RIS. We investigate how RIS-assisted systems can exhibit antifragile behavior under diverse and potentially severe jamming models, including Digital Radio Frequency Memory (DRFM) and phase/amplitude-shifting attacks, as well as passive and active RIS configurations. We adopt the RIS channel model used in several related papers, such as in [6, 7, 8]. The attackers’ channel models are modified and adapted from [4, 6, 9]. Recent RIS studies have addressed reactive jammers and emphasized the need for resilient systems. The paper [10] proposed a joint beamforming optimization scheme for active and passive transmission that maximizes both the secrecy rate and the system throughput, reducing the impact of the jamming attack. Contrary to this approach, we leverage the jammer as an information relaying node and therefore configure the RIS phases solely to boost the legitimate users SNR, with no mechanism to degrade the eavesdropping link. Another resilient RIS-assisted system is presented in [11]. This work proposes two partitioned RIS optimization schemes, which integrate beamforming with artificial noise injection to suppress the eavesdroppers SNR and improve the legitimate users secrecy capacity. In contrast to this approach, we harness the jammers signal as a constructive relay. Hence, our design refrains from artificial noise injection and instead leverages the jammers’ SNR to boost the achievable throughput in the system.
In sum, this paper provides a novel antifragile communication framework, which can however effectively integrate known techniques, such as beamforming and jammer piggybacking, with jammer localization and mitigation. To the best of our knowledge, this is the first work to analyze antifragility in RIS-assisted wireless communications.
III System Model
III-A Reference Scenarios
The system is illustrated in Figure 1. It includes a transmitter (source), receiver (destination), and one RIS node. Two types of jamming nodes are illustrated, one that eavesdrop either the source (Jammer 1) or the signal reflected by the RIS (Jammer 2). We refer to the Jammer 1 scenario as Source-Aware Eavesdropping and to the Jammer 2 scenario as RIS-Aware-Eavesdropping.
Figure 1: RIS-based jamming scenarios
III-B Channel Model
We consider a Rayleigh fading channel model for legitimate communication. The received signal at the destination node for a jammer free system is here given as [8, 6, 7],
@@ III-B Channel Model
Where $M$ denotes the number of elements in the legitimate RIS, $x(t)$ is the transmitted signal from the source to the destination node, $\mathbf{h}_{SR}=[h_{SR,1},..,h_{SR,k},...,h_{SR,M}]$ is the fading channel from the source to the RIS, with $h_{SR,k}=g_{SR,k}e^{-j\theta_{k}}\sqrt{(d_{SR})^{-\delta}}$ is the channel coefficient related to the element $k$, and $\mathbf{h}_{RD}=[h_{RD,1},...,h_{RD,l},...,h_{RD,M}]^{T}$ is the fading channel from the RIS to the destination, with $h_{RD,l}=g_{RD,l}e^{-j\psi_{k,l}}\sqrt{(d_{RD})^{-\delta}}$ is the channel coefficient related to the element $l$, $\delta$ denotes the path loss exponent, $g_{\mathrm{SR,k}}$ and $g_{\mathrm{RD,l}}$ the channel gains for S-to-R through element $k$ and R-to-D through element $l$, $d_{\mathrm{SR}}$ and $d_{\mathrm{RD}}$ the distance from the source to the RIS and RIS to the destination, respectively. Moreover, the correlation matrix $\mathbf{R}$ is defined as an $M\mathrm{x}M$ matrix, that defines the correlation coefficients $\rho_{i,j}$ between the $i^{th}$ and $j^{th}$ elements of the RIS. Thus, $\forall i,j,0\leq\rho_{i,j}\leq 1$ and $\rho_{i,j}=1$ for $i=j$. In addition, the reflection coefficients of the RIS are within the diagonal phase matrix $\Phi=\mathrm{diag}[e^{-j\phi_{1}},e^{-j\phi_{2}},....,e^{-j\phi_{M}}]$.
III-B1 Source-Aware Eavesdropping
Jammer 1 eavesdrops directly on the source node and then attacks the destination node. To build an antifragile strategy, it is necessary to model the channel in the presence of a jammer. As a jammer could have several types of behavior, it needs to be modeled accordingly.
In order to consider the impact of Jammer 1 on the received signal, we denote by $\mathbf{h}_{\mathrm{E1}}$ the eavesdropping CSI from the source to Jammer 1, and by $\mathbf{h}_{\mathrm{J1}}$ the jamming CSI from Jammer 1 to destination, which is given by $h_{\mathrm{s}}$ [4, 9]:
@@ III-B1 Source-Aware Eavesdropping
Where $\kappa$ is the Rician factor, $\sigma^{2}$ the average power, $L$ the number of paths, and $R_{s,i}$ the Rayleigh distributed amplitude and $\theta_{s,i}$ the uniformly distributed phase of the channel. When the Rician factor $\kappa=0$, the Rician fading channel coefficients are simplified to the Rayleigh fading coefficients. The received signal, considering the presence of a jammer, is given by:
@@
With $A$ being a multiplicative factor that can be defined based on the type of jamming attack (Table I), explained in Section III-C, and $\tau_{\mathrm{J1}}$ denotes the delay of the jamming path through Jammer 1.
III-B2 RIS-Aware Eavesdropping
We denote the eavesdropping CSI from the RIS to Jammer 2 by $\mathbf{h}_{\mathrm{RJ}}$ and by $\mathbf{h}_{\mathrm{J2}}$ the jamming CSI from Jammer 2 to destination, calculated using Eq. (2). The received signal $y(t)$ is a superposition of the legitimate signal defined by Eq. (1), and the jamming signal (Jammer 2), which is defined by $y_{J2}(t)$ and given as [6, 9],
@@ III-B2 RIS-Aware Eavesdropping
III-C Jammer Model
A reactive (repeater) jammer operates by capturing the waveform emitted by the legitimate transmitter, optionally applying a deterministic transformation, and retransmitting the modified signal with the objective of degrading the receiver’s performance. We adopt three categories of jammer models: digital radio frequency memory (DRFM), phase shifting (PS) and Amplitude shifting (AS), which are summarized in Table I. The DRFM attack retransmits the target signal on a sample-by-sample basis with a constant amplification gain denoted as $\beta_{a}$. Assuming that $h_{\mathrm{E1}}$ and $h_{\mathrm{J1}}$ are defined using Eq. (2), the DRFM jamming signal can be expressed as follows:
@@ III-C Jammer Model
Table I: Jammer models summarized [2].
@@ IV Antifragility Scheme Design
In order to achieve antifragile gain, several factors related to both the jammer and the legitimate signal must be considered. The system must first detect a jamming attack and estimate the jammers delay relative to the legitimate signal, $\tau_{\mathrm{J1}}$ or $\tau_{\mathrm{J2}}$. Additionally, the signals must be received orthogonally, either in time or space, to avoid overlap that would hinder demodulation and decoding. Thus, this section is organized as follows: first, we present general system assumptions about jamming detection and initial delay estimation. Then, the orthogonality strategy is discussed, leading to jamming classification, and finally, signal, modulation, and coding adaptation.
IV-A Jamming Detection and Delay Estimation
The jamming signal can be detected by evaluating the Bit Error Rate (BER) of the received signal. Without loss of generality, we assume that Reed-Solomon (RS) code is used; other codes would need to be separately handled. The presence of a jammer is detected when the BER exceeds a certain threshold, rendering the RS code unable to decode the message. When this condition occurs, a Cross-Correlation (CC) analysis is performed on the signal. The CC metric is obtained by correlating the received sequence $y$ with its time-aligned replica $\tilde{y}$; the presence of a peak is taken as evidence of a jammer. Because the delay estimate $\tau$ provides only coarse alignment-insufficient for symbol- or sample-level precision-we compute a complete cross-correlation (i.e., a sliding inner product) rather than a single dot product, given by [2]:
@@ IV-B Orthogonality Approach
For spatial separation, we estimate the angle of arrival (AoA). Because first-order multipath reflections dominate mmWave links, the AoA distribution exhibits sharp peaks [12]. We therefore apply a forward-backward spatially smoothed MUSIC algorithm [13]. The resulting AoAs for both desired and jamming signals enable beamforming-based spatial filtering at the receiver, implemented with a linearly constrained minimum-variance (LCMV) beamformer [14].
The beamforming weights $\mathbf{w}$ are computed using the covariance matrix $\mathbf{R}$, the constraint matrix and vector $\mathbf{C}$ and $\mathbf{f}$, respectively:
$\mathbf{w}=\mathbf{R}^{-1}\mathbf{C}\left(\mathbf{C}^{H}\mathbf{R}^{-1}\mathbf{C}\right)^{-1}\mathbf{f}$ (12)
The legitimate and malicious signal can be reconstructed using the following equation, i.e.,
$\mathbf{s}=\mathbf{w}^{H}\mathbf{X}$ (13)
If the receiver cannot separate the signals spatially, temporal orthogonality must be established.
For temporal orthogonality, the duration of the legitimate signal is reduced based on the estimated jamming delay, to prevent an overlap. Let $T_{\mathrm{active}}$ denote the time it takes for a signal to trigger the jamming process and $T_{\mathrm{OFF}}$ the time for the jamming process to stop. The legitimate signal must fulfill the property:
$T_{\mathrm{active}}\leq D(s_{\mathrm{legit}})<\tau_{\mathrm{j}}$ (14)
, with $D(\cdot)$ is the duration of the signal, and $\tau_{\mathrm{j}}$ the time it takes for the jamming signal $s_{\mathrm{legit}}$ to be received at the destination.
@@ IV-C Jamming Classification
For the classification of the DRFM jammer, we propose a stepwise classification process, which employs a novel similarity ratio normalizing cross-correlation by self-correlation (SC) to provide robust detection less sensitive to inherent signal structures. First, the self-correlation $R_{yy}(\tau)$ of the legitimate signal is calculated based on Eq. (10). Afterwards the SC maximum value is normalized along the signal duration, and used as reference, i.e.,
@@
Next, the cross-correlation between the jamming signal and the legitimate signal is calculated, given by Eq. (10). The maximum cross-correlation $CC_{max}$ value is also normalized.
@@
If classification confidence is insufficient, the identification routine is executed again. Once a jammer type is conclusively determined, the receiver logs the category, the estimated delay and other relevant metrics and then forwards this information to the transmitter. The transmitter subsequently initiates the waveform and coding-adaptation procedure. Malicious RIS devices can be similarly recognized by the propagation delay they impose: both passive and active malicious RIS configurations introduce latencies typically shorter than those incurred by signal-processing reactive jammers.
IV-D Signal Adaptation
Antifragile gains can only be obtained when the legitimate waveform is modified such that the jammer affects none of its information dimensions. Once such orthogonality is secured, i.e., via temporal, spatial, or spectral separation, the jammer’s emission can be exploited to convey additional information without corrupting the desired data. Signal adaptations therefore focus on preserving the legitimate signal’s phase, amplitude, and other dimensions, while treating the jamming waveform as an independent, utilizable resource. For an AS jammer that perturbs the signal envelope, the system remaps its waveform to an $M$-PSK constellation, thereby removing amplitude dependence. With a PS jammer, the transmitter adopts ASK whose constellation points reside exclusively in the positive real half-plane, allowing any $180^{\circ}$ phase inversions to be easily corrected. For a DRFM jammer no modulation change is necessary, since the strategy leverages the PSK signal’s coherent addition with the jammer’s delayed replica.
IV-E Modulation and Code Adaptation
Once the signal adaptation is completed, a new code rate is determined based on the increased SNR value. Therefore, we denote the additional jamming SNR as $\text{SNR}_{J}$ and legitimate SNR as $\text{SNR}_{L}$:

### v2

# A Cross-Layer Analysis of Network Antifragility with RIS-assisted Links under Jamming Attacks

@@ Abstract
Antifragility is an economics term defined as measure of (monetary) benefits gained from the adverse events and variability of the markets. This paper integrates for the first time the antifragility into the network based on communication links with Reconfigurable Intelligent Surface (RIS) affected by a jamming attack. We analyze whether antifragility can be achieved for several jamming models. Beyond the link-level gains, the results reveal how antifragile RIS-assisted links can be integrated into multi-hop systems to improve end-to-end network resilience, connectivity, and throughput under adversarial effects.
@@ I Introduction
Conventional resilient network systems aim to preserve or restore functionality under hostile conditions. By contrast, antifragility, as defined in economics, leverages uncertainty and adverse conditions to achieve performance gains[Taleb.2013]. In networking, antifragility has remained largely unexplored to date. With the emergence of multi-hop wireless links based on Reconfigurable Intelligent Surfaces (RIS), antifragility may become a salient feature in networks, due to the RIS property of using two-dimensional arrays of tunable reflecting elements to manipulate signal amplitude and phase, mitigating interference and extending coverage [9430683].
This paper proposes an analytical model of a multi-hop RIS-assisted network system, including a transmitter, a receiver, a RIS and 2 types of jammers, and derives antifragile performance gains of the network under jamming attacks. We analyze various jamming models on links, including Digital Radio Frequency Memory (DRFM) and phase and amplitude shifting. By leveraging the RIS reflections, we prove that the network can improve throughput under certain conditions, demonstrating a remarkably antifragile behavior. Specifically, when the baseline data rate is low and adversarial power is high, we observe a notable antifragile gain factor of up to $5\times$. Although the gain factor is smaller for larger baseline data rates, it still shows a notable performance gain. This work is the first one to integrate antifragitly of communication links into RIS enabled wireless mobile networks.
@@ II Related Work
Related work focuses on antifragility in communications, and our study is the first to explore antifragility in wireless networks with RIS in a cross-layer analysis. For link antifragility, paper [Lichtman.2017] presents a method to design frequency-shift keying (FSK) waveforms in order to exploit reactive jammers, effectively forcing an attacker, i.e., the jammer node,for performance gains of legitimate users. Building on this fundamental idea, the early work in [Lichtman.2018] studied the antifragile wireless links. More recently, [Ji.2023], basing its antifragility scheme on paper [Lichtman.2017] focused on reducing the outage probability in multi-relay networks. We extend the antifragility scope to consider various jamming models, signal separation and spoofing signal creation, which other works do not consider. Regarding the assumptions on link channel model, we adopt the known RIS channel model, e.g., from [9780032, bjornson2020rayleigh, Bensalem.11620241182024, 10086588]. The attacker’s channel model is adapted from [Lichtman.2017, 9780032]. Paper [Sun.2021] is related due to a joint beamforming optimization scheme that maximizes both the secrecy rate and the system throughput, reducing the impact of the jamming attack. By contrast, we model the jammer as an information relaying node and configure the RIS phases solely to boost the SNR of legitimate users, i.e., without mechanisms to degrade the eavesdropping link. Paper [tsiota2019jamming] offers a system-level view of jamming attacks, showing other network effects beyond link-level Tx-Rx-Jammer model. Paper [10066528] integrates beamforming with artificial noise injection to suppress the eavesdroppers. Instead, we assume the jammers signal as a constructive relay without artificial noise injection.
III Physical Layer Assumptions
The system model is illustrated in Figure 1, including source, receiver, one RIS node, and two jammers. Two types of jamming paths are illustrated, with Jammer 1 and Jammer 2 nodes. We refer to the path through Jammer 1 as Source-Aware Eavesdropping Path, and to Jammer 2 as RIS-Aware-Eavesdropping Path. To model the performance of the two paths in a cross layer analysis, it is necessary to start from the physical layer channel and jammer node models.
III-A Source- and RIS-aware eavesdropping paths model
For both models, the signal received at the destination node in a jammer free network system is given based on [Bensalem.11620241182024, 9780032, bjornson2020rayleigh],
@@ III-A Source- and RIS-aware eavesdropping paths model
Where $M$ denotes the number of RIS elements, $x(t)$ is the transmitted signal from the source to the destination node, $w(t)$ is the noise, $\mathbf{h}_{SR}=[h_{SR,1},..,h_{SR,k},...,h_{SR,M}]$ is the fading channel from the source to RIS, with $h_{SR,k}=g_{SR,k}e^{-j\theta_{k}}\sqrt{(d_{SR})^{-\delta}}$ is the channel coefficient related to the element $k$, and $\mathbf{h}_{RD}=[h_{RD,1},...,h_{RD,l},...,h_{RD,M}]^{T}$ is the fading channel from the RIS to the destination, with $h_{RD,l}=g_{RD,l}e^{-j\psi_{k,l}}\sqrt{(d_{RD})^{-\delta}}$ is the channel coefficient related to the element $l$, $\delta$ denotes the path loss exponent, $g_{\mathrm{SR,k}}$ and $g_{\mathrm{RD,l}}$ the channel gains for S-to-R through element $k$ and R-to-D through element $l$, $d_{\mathrm{SR}}$ and $d_{\mathrm{RD}}$ the distance from the source to the RIS and RIS to the destination, respectively. Moreover, the correlation matrix $\mathbf{R}$ is defined as an $M\mathrm{x}M$ matrix, that defines the correlation coefficients $\rho_{i,j}$ between the $i^{th}$ and $j^{th}$ elements of the RIS. Thus, $\forall i,j,0\leq\rho_{i,j}\leq 1$ and $\rho_{i,j}=1$ for $i=j$. In addition, the reflection coefficients of the RIS are within the diagonal phase matrix $\Phi=\mathrm{diag}[e^{-j\phi_{1}},e^{-j\phi_{2}},....,e^{-j\phi_{M}}]$.
Figure 1: Two hop network model with two jamming paths.
We denote by $\mathbf{h}_{\mathrm{E1}}$ the eavesdropping CSI from the source to Jammer 1, and by $\mathbf{h}_{\mathrm{J1}}$ the jamming CSI from Jammer 1 to destination, which is given by $h_{\mathrm{s}}$ [Lichtman.2017]:
@@
Where $\kappa$ is the Rician factor, $\sigma^{2}$ the average power, $L$ the number of paths, and $R_{s,i}$ the Rayleigh distributed amplitude and $\theta_{s,i}$ the uniformly distributed phase of the channel. When the Rician factor $\kappa=0$, the Rician fading channel coefficients are simplified to the Rayleigh fading coefficients. The received signal on the path with Jammer 1 is given by:
@@
With $A$ being a multiplicative factor that can be defined based on the type of jamming attack (Table I, Section III-B), and $\tau_{\mathrm{J1}}$ denotes the end-to-end path delay.
We denote the eavesdropping CSI from RIS to Jammer 2 by $\mathbf{h}_{\mathrm{RJ}}$ and by $\mathbf{h}_{\mathrm{J2}}$ the jamming CSI from Jammer 2 to the destination, calculated using Eq. (2). The received signal $y(t)$ is a superposition of the legitimate signal defined by Eq. (1), and the jamming signal (Jammer 2), which is defined by $y_{J2}(t)$ and taken from [9780032],
@@
III-B Jammer model
A reactive (repeater) jammer operates by capturing the signal emitted by the legitimate transmitter, optionally applying a deterministic transformation, and retransmitting the modified signal with the objective of degrading the receiver’s performance. We analyze three categories of jammer models: digital radio frequency memory (DRFM), phase shifting (PS) and Amplitude shifting (AS), as summarized in Table I. The DRFM attack retransmits the target signal on a sample-by-sample basis with a constant amplification gain denoted as $\beta_{a}$. Assuming that $h_{\mathrm{E1}}$ and $h_{\mathrm{J1}}$ are defined using Eq. (2), the DRFM jamming signal can be expressed as follows:
@@ III-B Jammer model
Table I: Jammer models summarized [Lichtman.2018].
@@ IV Antifragility Scheme Design
In this section, we first estimate the path delay and attack detection. To achieve antifragile gain, the network system must first detect a jamming attack and estimate the path delay with jammers relative to the legitimate path, $\tau_{\mathrm{J1}}$ or $\tau_{\mathrm{J2}}$. Additionally, the signals at the receiver must be received orthogonally, either in time or space, to avoid overlap that would hinder demodulation and decoding. We then provide a jamming classification, and based on signal adaptation, calculate the antifragile gain. The proposed antifragility gain model can be easilyt extended to network routing. For instance, by reporting antifragile gain metrics to the routing system, the system can dynamically update link weights or routing tables to favor links exhibiting performance gains under interference. In dense or distributed RIS deployments, this allows the network to exploit jamming-induced variability as a routing advantage, enabling self-optimizing, interference-aware topologies.
IV-A Jamming detection and path delay estimation
The jamming signal can be detected by evaluating the Bit Error Rate (BER) of the received signal. Without loss of generality, we assume that Reed-Solomon (RS) code is used. The presence of a jammer is detected when the received BER is such that the RS code is unable to decode the message. When this condition occurs, a Cross-Correlation (CC) analysis is performed on the signal. The CC metric is obtained by correlating the received sequence $y$ with its time-aligned replica $\tilde{y}$; the presence of a peak is taken as evidence of a jammer. Because the delay estimate $\tau$ provides only coarse alignment-insufficient for symbol- or sample-level precision-we compute a complete cross-correlation (i.e., a sliding inner product) rather than a single dot product, given by [Lichtman.2018]:
@@ IV-B Orthogonality Approach
For spatial separation, we estimate the angle of arrival (AoA). Because first-order multipath reflections dominate mmWave frequencies, the AoA distribution exhibits sharp peaks. We therefore apply a well known forward-backward spatially smoothed MUSIC algorithm based on [Pillai.1989]. The resulting AoAs for both desired and jamming signals enable beamforming-based spatial filtering at the receiver, implemented with a linearly constrained minimum-variance (LCMV) beamformer [Bourgeois.2009]. When the receiver cannot separate the signals spatially, temporal orthogonality is needed.
@@ IV-C Jamming classification
For the classification of the DRFM jammer, we propose a stepwise classification process, which employs a novel similarity ratio normalizing cross-correlation by self-correlation (SC) to provide robust detection less sensitive to inherent signal structures. This novel similarity ratio evaluates how the jamming signal resembles the legitimate signal. First, the self-correlation $R_{yy}(\tau)$ of the orthogonal legitimate signal is calculated based on Eq. (10). Afterwards the SC maximum value is normalized along the signal duration, and used as reference, i.e.,
@@
Next, the cross-correlation between the orthogonally received jamming signal $\hat{j}$ and the legitimate signal $\hat{y}$ is calculated, given by Eq. (10). The maximum cross-correlation $CC_{max}$ value is also normalized.
@@
IV-D Signal, modulation and code adaptation
Antifragile gains can only be obtained when the legitimate waveform is modified such that the jammer affects none of its information dimensions. Signal adaptations therefore focus on preserving the legitimate signal’s phase, amplitude, and other dimensions, while treating the jamming waveform as an independent, utilizable resource. For an AS jammer that perturbs the signal envelope, the system remaps its waveform to an $M$-PSK constellation, thereby removing amplitude dependence. With a PS jammer, the transmitter adopts ASK whose constellation points reside exclusively in the positive real half-plane, allowing any $180^{\circ}$ phase inversions to be easily corrected. For a DRFM jammer no modulation change is necessary, since the strategy leverages the PSK signal’s coherent addition with the jammer’s delayed replica.
Once the signal is adapted, a new code rate is determined based on the increased SNR value. Therefore, we denote the jamming SNR as $\text{SNR}_{J}$ and legitimate SNR as $\text{SNR}_{L}$: